# Likelihood Based Inference for ARMA Models

### v5

@@ 31                         Set $\hat{\psi}=\operatorname*{arg\,max}_{j\in 0:k}\ell\big{(}\hat{\psi}_{j}\big{)}$;
Algorithm 1 MLE for ARMA Models.
@@ 2 Inputs (defaults):
5                                                                                                                                                         Probability of sampling a root pairs as real $0\leq p\leq 1$, ($p=\sqrt{1/2}$).
@@
7                                                                                                                                                         Numeric optimization routine $f(\psi)$ (Gardner et al., 1980).
@@ 2.2 Simulation Studies
To investigate the extent to which the standard approach for ARMA parameter estimation results in improperly maximized likelihoods, we conduct a series of simulation studies. It is challenging to obtain precise estimates of how frequently current standards lead to sub-optimal parameter estimates due to the varied applications of ARMA models in practice, the diversity in data sizes ($n$) and model orders $(p,q)$, and the differing degrees to which an ARMA model adequately describes the data-generating process. Therefore, we restrict our simulation studies to idealized scenarios where the data-generating process is Gaussian-ARMA, recognizing that likelihood maximization is easiest for this model class, thereby resulting in conservative estimates.
@@
In the first simulation study, we simulate time series data of lengths $n\in\{50,100,500,1000\}$ from Gaussian-ARMA models with known orders $(p,q)\in\{1,2,3\}^{2}$, generating 36,000 unique models and datasets. We avoid any models that contain parameter redundancies by requiring the data generating model to have a minimum distance of $0.1$ between all roots of $\Phi(x)$ and $\Theta(x)$. We further restrict model coefficients so that they do not lie near boundary conditions. Models of the same order of the generating data are fit to the data, simplifying the the problem further by avoiding the order selection step that is necessary in most data analyses. In doing so, we attempt to answer the question of how often sub-optimal estimates may arise using existing software in the case where the parameter estimation procedure should be as easy as possible for the given combinations of $(n,p,q)$.
@@
Even in this extremely simplified scenario, existing software failed to properly maximize model likelihoods in at least $20.8\%$ of the simulated datasets—evidenced by an improvement obtained using Algorithm 1. Though this improvement may appear modest, an improvement in $20.8\%$ of the large number of published ARMA models would affect many papers. Furthermore, time series analysis courses and textbooks often recommend fitting multiple ARMA models to a dataset. Consequently, the probability that at least one candidate model is not properly optimized increases significantly in practice. The rate of improvement obtained using our algorithm increases with model complexity and decreases with more observations (Fig. 2). For example, likelihoods improved in $55.1\%$ of the simulations when $p=q=3$ and $n=50$; models of this size and number of observations are not uncommon in published research studies.
Importantly, we do not claim to improve likelihood for all ARMA models and datasets. In this simulation study, our likelihood maximum routine did not improve likelihoods for many of the simulated data. However, these results demonstrate that there is a very real potential for obtaining sub-optimal parameter estimates when using only a single parameter initialization, even in the most idealistic scenarios. Rather than having to worry if a single initialization is sufficient to fit a given model, it is preferable to adopt methods that make such situations rare. The primary limitation of our algorithm is that the potential for improved fits comes at the cost of increased computation times. In our simulation study, however, the average time to estimate parameters using our approach was $0.6$ seconds, a computational expense that is worth the effort in many situations.
@@
Figure 2: Proportion of simulated data with improved likelihood from using multiple restarts (Algorithm 1).
@@
The median log-likelihood improvement in this simulation study was $0.66$, with an interquartile range of $(0.22,1.46)$. Among the most common motivations for fitting ARMA models is to model serial correlations in a regression model; in this setting, the discovered shortcomings in log-likelihood are often enough to change the outcome of the analysis. For instance, consider modeling $y_{i}=\beta x_{i}+\epsilon_{i}$, where $\beta\in\mathbb{R}$, and we model the error terms $\epsilon_{i}\sim\text{ARMA}(p,q)$. For now, we will assume the order $(p,q)$ is fixed. We may wish to test the hypothesis $H_{0}:\beta=0$ vs $H_{1}:\beta\neq 0$. A standard approach to doing this is a likelihood ratio test, and using Wilks’ theorem to get an approximate test. We denote $ll_{0}$ and $ll_{1}$ as the maximum log-likelihood of the model under $H_{0}$ and $H_{1}$, respectively. The standard approximation is to assume $2\Delta=2(ll_{1}-ll_{0})\sim\chi^{2}_{1}$. Using a significance level of $\alpha=0.05$, we would reject $H_{0}$ if $\Delta\geq 1.92$. Given that $E_{H_{0}}[\Delta]=0.5$, subtracting the reported log-likelihood deficiencies (which has a median value of $0.66$) of existing software to either or both $ll_{0},ll_{1}$ could change the outcome of this test.
@@ 2.2.1 Parameter Uncertainty
Improved parameter estimation leads to modified standard error estimates, which are default outputs in R and Python. These standard errors result from the numeric optimizer’s estimate of the gradient of the log-likelihood, used to approximate the Fisher information matrix. These standard errors, though not inherently of interest, are sometimes used justify the inclusion of a parameter in a model (Brockwell and Davis, 1991, Chatper 9). We extend our simulation study to examine this approach and how our algorithm impacts the estimates. For each of the 36,000 generative models from the previous study, we generate 100 additional datasets, estimating the MLE and Bonferroni-adjusted $95\%$ confidence intervals using both estimated standard errors and profile likelihood confidence intervals (PLCIs) from Wilks’ theorem. Figure 3 shows PLCIs had better or equivalent nominal coverage than Fisher-based confidence intervals across all combinations of $p$, $q$, and $n$. We found that the interval estimation method mattered more for confidence interval performance than the specific parameter estimation algorithm. The relevance of this result is explored further in Section 3.1.
Figure 3: Proportion of models that achieved nominal coverage of Bonferroni adjusted $95\%$ confidence intervals. The dashed line denotes the target coverage level. (A) Confidence intervals created using Fisher’s information matrix. (B) Confidence intervals created using profile likelihoods.
@@ 2.2.2 AIC Table Consistency
A common contemporary approach involves fitting several candidate models and selecting the one that minimizes a criterion like Akaike’s information criterion (AIC) (Akaike, 1974). This can be done by explicitly creating a table of all candidate models and their corresponding AIC values; in this case issues of improper maximization become more apparent. For instance, a table of AIC values may contain numeric inconsistencies, where a larger model may have lower estimated likelihoods than a smaller model within which it is nested (for an example, see Section 3). This type of result can make a careful practitioner feel uneasy, as there is evidence that at least one candidate model was not properly optimized. Evidence of improper optimization may be less evident when relying on software that automates this process, such as the automatic procedure of Hyndman and Khandakar (2008), but the potential for sub-optimal estimates remains.
@@
We conducted an additional simulation study to investigate numeric inconsistencies that may arise when fitting multiple model parameters. As before, we simulated 1000 unique models and datasets of size $n\in\{50,100,500,1000\}$ from Gaussian ARMA$(p,q)$ models for $(p,q)\in\{1,2,3\}^{2}$. To avoid models with parameter redundancies, we ensured a minimum distance of 0.1 between all roots of $\Phi(x)$ and $\Theta(x)$ and excluded models with coefficients near boundary conditions. For each dataset, AIC tables were created for model sizes $(p,q)\in\{0,1,2,3\}^{2}$.
@@
The single parameter initialization approach resulted in AIC table inconsistencies in $45.6\%$ of the simulated datasets. Although our proposed algorithm significantly mitigates this issue, it does not guarantee that all model likelihoods are fully maximized. This is illustrated in Figure 4, where a non-zero percentage of AIC tables remain inconsistent, even as the algorithm’s stopping criterion grows. The ARMA(1,1)11(1,1)( 1 , 1 ) panel in Figure 4 illustrates the increasing difficulty of parameter estimation when dealing with parameter redundancies. In such cases, it is often necessary to adjust additional parameters in the numeric optimization routine. For example, our R implementation of the algorithm relies on the generic BFGS optimizer in the stats::optim function. Modifying the optimization method or the default hyperparameters can lead to improved fits or faster convergence rates. While the default parameters of the numeric optimizer are generally adequate, increasing the maximum number of algorithmic iterations can be beneficial.
The contemporary approach of using AIC—or any other information based criteria—to select model order involves fitting unnecessarily large models, leading to parameter redundancies that complicate likelihood optimization. The use of AIC for ARMA model order selection has theoretical support, particularly for forecasting, as ARMA models inspired the original AIC paper (Akaike, 1974). However, without proper likelihood maximization, a strategy that considers only a single parameter initialization may not truly minimize AIC. In this framework, likelihood maximization and over-parameterization are interconnected: all candidate models must be maximized for likelihood, or users risk selecting over-parameterized models that fail to minimize the intended information criterion.
Classical ARMA modeling addresses this by recommending diagnostic plots to determine appropriate model order and advising against simultaneously adding AR and MA components. In this approach, the additional difficulty in parameter estimation associated with fitting models containing parameter redundancies is avoided by not fitting overly complex models when possible. Despite this, shortcomings in likelihood maximization can occur even in models without parameter redundancies (Figs. 12, and 4), necessitating the exploration of multiple parameter initializations.
@@
Figure 4: Data is generated from ARMA$(p,q)$ models with $(p,q)\in\{1,2,3\}^{2}$, and the corresponding AIC table is created. The Y-axis shows the percentage of tables that were consistent. M is the number of times a maxima is observed before the algorithm terminates, so $M=1$ corresponds to the standard maximization procedure.
3 Annual Depths of Lake Michigan

### v6

@@ 30          Set $\hat{\psi}=\operatorname*{arg\,max}_{j\in 0:k}\ell\big(\hat{\psi}_{j}\big)$;
Algorithm 1 MLE with Uniform-Root Sampling.
@@ 2 Inputs (defaults):
6                                    Numeric optimization routine $f(\psi)$ ([3]).
@@
For the first approach, it is an easier task to sample inverted roots, as the sufficient conditions for causality and invertibility now require that the inverted roots lie inside the complex unit circle, a region easier to sample uniformly. The roots of the polynomials can be real or complex; complex roots must come in complex conjugate pairs in order for all of the corresponding model parameters to be real. The simplest approach would be to sample all inverted root pairs ($z_{1},z_{2}$) within the complex unit circle by uniformly sampling angles and radii (lines 9-14 of Algorithm 1). However, this would imply almost surely all root pairs are complex, and some model parameters would only be sampled as positive (or negative). For instance, consider an AR(2) model. The AR polynomial is:
In this equation, if both $z_{1},z_{2}$ are complex conjugates, then $\phi_{2}=z_{1}z_{2}>0$. As such, the only way that $\phi_{2}<0$ is if $z_{1},z_{2}\in\mathbb{R}$. Similar results hold for the MA coefficients with opposite signs for the coefficients. This issue is directly addressed in lines 5-8 of Algorithm 1: root pairs are sampled as real with probability $p=\sqrt{1/2}$, and real pairs sampled with the same sign with probability $p$, such that the product (and sums) of each pair is positive with probability $1/2$. We sample conjugate pairs within an annular disk on the complex plane to avoid trivial and approximately non-stationary cases. The radii of both the inner and outer circles defining the disk are defined using $\gamma$ in lines 7, 10, and 13 of Algorithm 1.
Algorithm 1 is designed to directly sample the roots of the AR and MA polynomials, as the roots determine the causality and invertibility of the resulting model. However, it is not immediately evident how the choice of sampling a proportion $p=\sqrt{1/2}$ will effect the final distribution of ARMA coefficients, other than ensuring that parameters are sampled as positive or negative with similar frequencies. Existing theory can be used to provide an alternative approach to sampling model parameters within the admissibility region [18]. Algorithm 2 leverages existing techniques by first generating a vector of partial autocorrelations, each taking values in the interval (1,1)(-1,1). These vectors are transformed into ARMA parameters using the Durbin-Levinson recursion, ensuring the resulting parameters are within the admissible region. This process is outlined in equations (4) and (5) of [10], who demonstrate that the established theory for this transformation continues to hold at boundary conditions.
6Get preliminary estimate: $\hat{\psi}_{0}=f(\psi_{0})$; set $k=0$;
7 repeat Until stopping criterion met
8   Increment $k$;
9   if $p$ (or $q$) =1=1 then
10     Sample $\phi_{1}$ (or $\theta_{1}$) uniformly from $U(-1+\gamma,1-\gamma)$;
11     Set corresponding root $z^{\text{AR}}_{1}$ (or $z^{\text{MA}}_{1}$) to be $\phi_{1}$ (or $\theta_{1}$);
12
13     else
14       Set AR and MA roots $\{z_{i}^{\text{AR}}\}_{i=1}^{p}=0_{p}$, $\{z_{j}^{\text{MA}}\}_{j=1}^{q}=0_{q}$;
15
16       Increment $k$;
17       while $\min_{i,j}|z^{\text{AR}}_{i}-z^{\text{MA}}_{j}|<\alpha$ do
18         Uniformly sample $\xi^{\text{AR}}_{i,i},\,\xi^{\text{MA}}_{j,j}\overset{i.i.d.}{\sim}U(\gamma,1-\gamma)$ for $i\in 1\!:\!p,\,j\in 1\!:\!q$;
19         for $i$ in 2:p2\!:\!p do
20           for $h$ in 1:(i1)1\!:\!(i-1) do
21             $\xi^{\text{AR}}_{i,h}=\xi^{\text{AR}}_{i-1,h}-\xi^{\text{AR}}_{i,i}\xi^{\text{AR}}_{i-1,i-h}$;
22
23
24             for $j$ in 2:q2\!:\!q do
25               for $h$ in 1:(j1)1\!:\!(j-1) do
26                 $\xi^{\text{MA}}_{j,h}=\xi^{\text{MA}}_{j-1,h}-\xi^{\text{MA}}_{j,j}\xi^{\text{MA}}_{j-1,j-h}$;
27
28
29                 Set $\psi_{k}=(\phi^{k}_{1},\ldots,\phi^{k}_{p},\theta^{k}_{1},\ldots,\theta^{k}_{q})=(\xi^{\text{AR}}_{p,1},\ldots,\xi^{\text{AR}}_{p,p},-\xi^{\text{MA}}_{q,1},\ldots,-\xi^{\text{MA}}_{q,q})$;
30                 Calculate roots $\{z_{j}^{\text{AR}}\}_{j=1}^{p}$, $\{z_{j}^{\text{MA}}\}_{j=1}^{q}$ from sampled coefficients;
31
32                 Estimate $\hat{\psi}_{k}=f(\psi_{k})$;
33
34                 until ;
Algorithm 2 MLE with Durbin-Levinson Sampling.
2 Inputs (defaults):
5                                    Numeric optimization routine $f(\psi)$ ([3]).
Parameter redundancy in ARMA models occurs when the polynomials $\Phi(x)$ and $\Theta(x)$ share one or more roots, leading to an overall reduction in model order. This complicates parameter initialization, optimization, and identifiability. The ARMA model (Eq. 1) can be rewritten as:
$\Phi(B)x_{t}=\Theta(B)w_{t},$ (5)
where $B$ is the backshift operator, i.e., $Bx_{t}=x_{t-1}$. Using the fundamental theorem of algebra, Equation 5 can be factored into
$(1-\lambda_{1}B)\ldots(1-\lambda_{p}B)x_{t}=(1-\nu_{1}B)\ldots(1-\nu_{q}B)w_{t},$
where $\{\lambda_{i}\}_{i=1}^{p}$ and $\{\nu_{j}\}_{j=1}^{q}$ are the inverted roots of $\Phi(B)$ and $\Theta(B)$, respectively. If $\lambda_{i}=\nu_{j}$ for any $(i,j)\in\{1,\ldots,p\}\times\{1,\ldots,q\}$, then the roots will cancel each other out, resulting in an ARMA model of smaller order. As an elementary example, consider the $\text{ARMA}(1,1)$ and $\text{ARMA}(2,2)$ models in equations 6 and 7.
$x_{t}$ $=\frac{1}{3}x_{t-1}+w_{t}+\frac{2}{3}w_{t-1},$ (6)
$x_{t}$ $=\frac{5}{6}x_{t-1}-\frac{1}{6}x_{t-2}+w_{t}+\frac{1}{6}w_{t-1}-\frac{1}{3}w_{t-2}.$ (7)
While these two models appear distinct at first glance, re-writing the models in polynomial form (Eq. 5) shows that these two models are actually equivalent after canceling out the common factors on each side of the equation.
In a similar fashion, it is possible that the roots are not exactly equal but are approximately equal. In this case, the ratio of factors becomes close to one, resulting in a similar effect to when the roots exactly cancel. In Algorithms 1 and 2, we avoid the possibility of nearly canceling roots in parameter initializations by requiring the minimum Euclidean distance between inverted polynomial roots to be greater than $\alpha$. This is done in line 4 of Algorithm 1 and line 10 of Algorithm 2, though the condition is rarely triggered if the order of the model is of typical size $(p,q<4)$.
Both of the sampling schemes are combined with existing procedures for numeric optimization of model log-likelihoods (lines 16 and 20, respectively), as well as the default initialization strategy for $\psi_{0}$ used by existing software (such as the CSS initialization, in line 1 of both algorithms). Doing so guarantees that final estimates obtained from both algorithms correspond to likelihood values greater than or equal to the currently accepted standards in the software environment where the algorithms are implemented. For this article, both the numeric optimization procedure $f(\cdot)$ and the parameter initialization strategy to obtain $\psi_{0}$ are those implemented in stats::arima. The stopping criterion was chosen so that the algorithm stops trying new initial values when no new maximum has been found using the last $M$ parameter initializations. Alternative stopping criterion can be used (see for example, [19]), but we found that this simple heuristic works well in practice.
Both algorithms are implemented in the R package arima2, available on the Comprehensive R Archive Network (CRAN) [20]. The package features the function arima2::arima, which is an adaptation of the stats::arima function modified to incorporate the new initialization schemes. The choice of initialization scheme can be selected using the init_method argument to the function.
3 Results
@@ 3.1 Simulation Studies
To investigate the extent to which the standard approach for ARMA parameter estimation results in improperly maximized likelihoods, we conduct a series of simulation studies. It is challenging to obtain precise estimates of how frequently current standards lead to sub-optimal parameter estimates due to the varied applications of ARMA models in practice, the diversity in data sizes ($n$) and model orders $(p,q)$, and the differing degrees to which an ARMA model adequately describes the data-generating process. Therefore, we restrict our simulation studies to idealized scenarios where the data-generating process is Gaussian-ARMA, recognizing that likelihood maximization is easiest for this model class, thereby resulting in conservative estimates of how frequently our algorithm improves model likelihood.
@@
In the first simulation study, we simulate time series data of lengths $n\in\{50,100,500,1000\}$ from Gaussian-ARMA models with known orders $(p,q)\in\{1,2,3\}^{2}$. For each combination of (p, q, n), 1000 unique parameter sets are generated, as well as a single time series from each set of parameters. In total, this results in 36,000 unique models and datasets.We avoid any models that contain parameter redundancies by requiring the data generating model to have a minimum distance of $0.1$ between all roots of $\Phi(x)$ and $\Theta(x)$. We further restrict model coefficients so that they do not lie near boundary conditions. Models of the same order of the generating data are fit to the data, simplifying the the problem further by avoiding the order selection step that is necessary in most data analyses. In doing so, we attempt to answer the question of how often sub-optimal estimates may arise using existing software in the case where the parameter estimation procedure should be as easy as possible for the given combinations of $(n,p,q)$.
@@
Even in this extremely simplified scenario, existing software failed to properly maximize model likelihoods in at least $23.4\%$ of the simulated datasets—evidenced by an improvement obtained using either Algorithm 1 or Algorithm 2. Though this improvement may appear modest, an improvement in $23.4\%$ of the large number of published ARMA models would affect many papers—a number measured in thousands of papers since 2024 following our estimate in the introduction. Furthermore, time series analysis courses and textbooks often recommend fitting multiple ARMA models to a dataset, and here we only fit one for each algorithm. Consequently, the probability that at least one candidate model is not properly optimized increases significantly in practice. The rate of improvement obtained using our algorithm increases with model complexity and decreases with more observations (Fig. 2). For example, likelihoods improved in $61.9\%$ of the simulations when $p=q=3$ and $n=50$ by using both algorithms. Models of this size and number of observations are not uncommon in published research studies.
In this simulation study, we found that Algorithm 1 generally resulted in greater improvements over the baseline compared to Algorithm 2. Specifically, Algorithm 1 improved the log-likelihood in $20.7\%$ of the simulated data relative to the baseline, whereas Algorithm 2 did so in $20.4\%$ of cases. Moreover, when the two algorithms produced different outputs, Algorithm 1 more often achieved the highest log-likelihood. For simplicity, the remainder of the article summarizes the results obtained using only Algorithm 1 rather than both algorithms. However, the fact that some improvements were achieved by Algorithm 2 alone indicates that it may be more suitable for certain datasets, or that increasing the number of random initializations considered by having a more strict stopping criteria may be beneficial.
@@
Fig 2: Proportion of simulated data with improved likelihood from using multiple restarts (Algorithm 1).
Importantly, we do not claim to improve likelihood for all ARMA models and datasets. In this simulation study, our likelihood maximum routine did not improve likelihoods for many of the simulated data. However, these results demonstrate that there is a very real potential for obtaining sub-optimal parameter estimates when using only a single parameter initialization, even in the most idealistic scenarios. Rather than having to worry if a single initialization is sufficient to fit a given model, it is preferable to adopt methods that make such situations rare. The primary limitation of our algorithm is that the potential for improved fits comes at the cost of increased computation times. In our simulation study, however, the average time to estimate parameters using our approach was $0.6$ seconds, a computational expense that is worth the effort in many situations. Average computing times from our simulation study are given in Table 1. The amount of computing time required for each time series depends heavily on many factors, including the proximity of parameter initializations to local optimums. When $n$ is large, the CSS initialization will be closer to one of these local optimums, resulting in much faster convergence than randomly initialized parameters.
Table 1: Table summarizing the computing times of each simulated dataset, in seconds. For large $n$, the CSS initialization is generally close to a maximum, causing the default initialization to converge quicker than random initializations.
(a) Algorithm 1.
$n$ Mean SD
50 0.28 0.20
100 0.33 0.24
500 0.74 0.58
1000 1.13 0.90
(b) Algorithm 2.
$n$ Mean SD
50 0.30 0.21
100 0.35 0.25
500 0.78 0.60
1000 1.20 0.92
(c) stats::arima.
$n$ Mean SD
50 0.03 0.02
100 0.03 0.02
500 0.05 0.04
1000 0.07 0.06
@@
The median log-likelihood improvement in this simulation study was $0.66$, with an interquartile range of $(0.22,1.47)$. Among the most common motivations for fitting ARMA models is to model serial correlations in a regression model; in this setting, the discovered shortcomings in log-likelihood are often enough to change the outcome of the analysis. For instance, consider modeling $y_{i}=\beta x_{i}+\epsilon_{i}$, where $\beta\in\mathbb{R}$, and we model the error terms $\epsilon_{i}\sim\text{ARMA}(p,q)$. For now, we will assume the order $(p,q)$ is fixed. We may wish to test the hypothesis $H_{0}:\beta=0$ vs $H_{1}:\beta\neq 0$. A standard approach to doing this is a likelihood ratio test, and using Wilks’ theorem to get an approximate test. We denote $ll_{0}$ and $ll_{1}$ as the maximum log-likelihood of the model under $H_{0}$ and $H_{1}$, respectively. The standard approximation is to assume $2\Delta=2(ll_{1}-ll_{0})\sim\chi^{2}_{1}$. Using a significance level of $\alpha=0.05$, we would reject $H_{0}$ if $\Delta\geq 1.92$. Given that $E_{H_{0}}[\Delta]=0.5$, subtracting the reported log-likelihood deficiencies (which has a median value of $0.66$) of existing software to either or both $ll_{0},ll_{1}$ could change the outcome of this test.
Likelihood optimization for mixed ARMA models is known to be more challenging than for pure AR or MA processes. Nonetheless, Figure 1 demonstrates that multimodal likelihood surfaces and suboptimal estimates can arise in pure MA(1) processes, and even these simple models can benefit from multiple initializations. This class of models is particularly useful for illustrating the potential for suboptimal estimates, as the presence of only a single parameter allows for easy visualization. Our preliminary results indicate, however, that pure AR and MA processes rarely require multiple initializations for effective optimization. The difficulty of parameter optimization in mixed ARMA models is one reason authors have previously advised caution in their use [21]. Despite the need for caution, mixed models remain commonly used in practice and are therefore the primary focus of our article.
@@ 3.1.1 Parameter Uncertainty
Improved parameter estimation leads to modified standard error estimates, which are default outputs in R and Python. These standard errors result from the numeric optimizer’s estimate of the gradient of the log-likelihood, used to approximate the Fisher information matrix. These standard errors, though not inherently of interest, are sometimes used justify the inclusion of a parameter in a model [22, Chatper 9]. We extend our simulation study to examine this approach and how our algorithm impacts the estimates. For each of the 36,000 generative models from the previous study, we generate 100 additional datasets, estimating the MLE and Bonferroni-adjusted $95\%$ confidence intervals using both estimated standard errors and profile likelihood confidence intervals (PLCIs) from Wilks’ theorem. The accuracy of confidence intervals is evaluated by comparing the nominal coverage of the confidence intervals from the simulations to the target $95\%$ coverage level.  Fig 3 shows PLCIs had better or equivalent nominal coverage than Fisher-based confidence intervals across all combinations of $p$, $q$, and $n$. We found that the interval estimation method mattered more for confidence interval performance than the specific parameter estimation algorithm. The relevance of this result is explored further in Section 3.2.1.
Fig 3: Proportion of models that achieved nominal coverage of Bonferroni adjusted $95\%$ confidence intervals. The dashed line denotes the target coverage level. Parameter estimates were obtained using Algorithm 1. (A) Confidence intervals created using Fisher’s information matrix. (B) Confidence intervals created using profile likelihoods.
@@ 3.1.2 AIC Table Consistency
A contemporary approach involves fitting several candidate models and selecting the one that minimizes a criterion like Akaike’s information criterion (AIC) [23]. This can be done by explicitly creating a table of all candidate models and their corresponding AIC values; in this case issues of improper maximization become more apparent. For instance, a table of AIC values may contain numeric inconsistencies, where a larger model may have lower estimated likelihoods than a smaller model within which it is nested (for an example, see Section 3.2). This type of result can make a careful practitioner feel uneasy, as there is evidence that at least one candidate model was not properly optimized. Evidence of improper optimization may be less evident when relying on software that automates this process, such as the automated Hyndman-Khandakar algorithm [24], but the potential for sub-optimal estimates remains.
@@
We conducted an additional simulation study to investigate numeric inconsistencies that may arise when fitting multiple model parameters. As before, we simulated 1000 unique models and datasets of size $n\in\{50,100,500,1000\}$ from Gaussian ARMA(p,q)(p,q) models for $(p,q)\in\{1,2,3\}^{2}$. To avoid models with parameter redundancies, we ensured a minimum distance of 0.1 between all roots of $\Phi(x)$ and $\Theta(x)$ and excluded models with coefficients near boundary conditions. For each dataset, AIC tables were created for model sizes $(p,q)\in\{0,1,2,3\}^{2}$.
@@
The single parameter initialization approach resulted in AIC table inconsistencies in $45.6\%$ of the simulated datasets. Although our proposed algorithm significantly mitigates this issue, it does not guarantee that all model likelihoods are fully maximized. This is illustrated in Fig 4, where a non-zero percentage of AIC tables remain inconsistent, even as the algorithm’s stopping criterion grows. The ARMA(1,1)(1,1) panel in Fig 4 illustrates the increasing difficulty of parameter estimation when dealing with parameter redundancies. In such cases, it is often necessary to adjust additional parameters in the numeric optimization routine. For example, our R implementation of the algorithm relies on the generic BFGS optimizer in the stats::optim function. Modifying the optimization method or the default hyperparameters can lead to improved fits or faster convergence rates. While the default parameters of the numeric optimizer are generally adequate, increasing the maximum number of algorithmic iterations can be beneficial for fully maximizing the likelihood for challenging models and data.
@@
Fig 4: Data is generated from ARMA(p,q)(p,q) models with $(p,q)\in\{1,2,3\}^{2}$, and the corresponding AIC table is created. The Y-axis shows the percentage of tables that were consistent. M is the number of times a maxima is observed before the algorithm terminates, so $M=1$ corresponds to the standard maximization procedure.
The contemporary approach of using AIC—or any other information based criteria—to select model order involves fitting unnecessarily large models, leading to parameter redundancies that complicate likelihood optimization. The use of AIC for ARMA model order selection has theoretical support, particularly for forecasting, as ARMA models inspired the original AIC paper [23]. However, without proper likelihood maximization, a strategy that considers only a single parameter initialization may not truly minimize AIC. In this framework, likelihood maximization and over-parameterization are interconnected: all candidate models must be maximized for likelihood, or users risk selecting over-parameterized models that fail to minimize the intended information criterion. For the current study, the choice of AIC versus other popular information criteria such as the corrected AIC (AICC) or Bayesian information criterion (BIC) is unimportant: all of these approaches rely on proper optimization of the likelihood function, which is the problem we are addressing here.
Classical ARMA modeling addresses this by recommending diagnostic plots to determine appropriate model order and advising against simultaneously adding AR and MA components. In this approach, the additional difficulty in parameter estimation associated with fitting models containing parameter redundancies is avoided by not fitting overly complex models when possible. Despite this, shortcomings in likelihood maximization can occur even in models without parameter redundancies (Figs 12, and 4), necessitating the exploration of multiple parameter initializations. Further, the increasing preference of using automated software to pick the model size using an information criterion suggests the importance of using software that reliably maximizes model likelihoods even in the presence of over-parameterization.
3.2 Annual Depths of Lake Michigan